# Never too Cocky to Cooperate: An FIM and RL-based USV-AUV Collaborative System for Underwater Tasks in Extreme Sea Conditions

### v1

@@ Abstract
This paper develops a novel unmanned surface vehicle (USV)–autonomous underwater vehicle (AUV) collaborative system designed to enhance underwater task performance in extreme sea conditions. The system integrates a dual strategy: (1) high-precision multi-AUV localization enabled by Fisher information matrix-optimized USV path planning, and (2) reinforcement learning-based cooperative planning and control method for multi-AUV task execution. Extensive experimental evaluations in the underwater data collection task demonstrate the system’s operational feasibility, with quantitative results showing significant performance improvements over baseline methods. The proposed system exhibits robust coordination capabilities between USV and AUVs while maintaining stability in extreme sea conditions. To facilitate reproducibility and community advancement, we provide an open-source simulation toolkit available at: https://github.com/360ZMEM/USV-AUV-colab .
@@ I Introduction
Despite recent advances, existing USV-AUV frameworks often lack the intelligence and adaptability required for complex tasks in extreme sea conditions [10]. For instance, while Ultra-Short Baseline (USBL) positioning via USV improves localization, most approaches either fix the USV at static coordinates or employ heuristic path planning [11]. Such rigidity results in inefficient USV-AUV linkage and fails to adapt to dynamic ocean environments, leading to suboptimal positioning and task performance. Further, traditional AUV control strategies (e.g., PID, fuzzy control) rely heavily on precise dynamic models, exhibit poor adaptability to turbulence, and struggle with multi-objective optimization or high-dimensional sensor data [12]. Crucially, many studies treat ocean dynamics as negligible noise—an unrealistic assumption in extreme conditions [13].
@@
To address these challenges, a dual strategy is essential: adaptive planning and control policy to enable dynamic decision-making in uncertain environments without relying on rigid models, while optimizing multi-AUV collaboration for efficient task execution, and uncertainty and information-aware path planning method to improve the accuracy of AUV positioning via USV to enhance the robustness of the USV-AUV system, mitigating localization errors caused by extreme sea conditions.
@@
This motivates the integration of reinforcement learning (RL) and Fisher Information Matrix (FIM) optimization—a synergistic approach that jointly tackles control and localization challenges [14]. On one hand, RL’s model-free adaptive optimization enables AUVs to learn robust policies through environmental interaction, overcoming nonlinearities and time-varying disturbances without precise dynamic models [15, 16]. Multi-Agent RL (MARL) further enhances cooperative planning and control, enabling efficient multi-AUV coordination under communication constraints [17]. On the other hand, FIM-based optimization directly quantifies and minimizes positioning uncertainty via USV path planning for maximizing the system’s information content, thereby reducing the Cramér-Rao Bound (CRB) for AUV navigation [14]. By combining RL’s adaptability with FIM’s precision, the proposed system achieves both intelligent decision-making and high-accuracy localization—a critical requirement for extreme sea scenarios.
@@
Based on above analysis, this paper develops a FIM and RL-based USV-AUV collaborative system for underwater tasks in extreme sea conditions to improve the mission complete efficiency, and contributions are listed as follows:
@@
  • A novel USV-AUV collaborative system is developed, which integrates a FIM and RL-based dual strategy to enable dynamic decision-making and uncertainty and information-aware capability to improve the mission complete efficiency for underwater tasks.
@@
  • This system realizes accurate positioning of AUVs via USV path planning through minimizing the determinant of the system’s FIM. Based on this, through integrating environment-aware ability into state space, and shaped reward function for multi-AUV collaboration, we further use RL to empower multi-AUV with intelligence and the adaptability to the underwater task.
@@
  • Through comprehensive experiments in the underwater data collection task, our system showcases superior feasibility and excellent performance with robustness in extreme sea conditions.
@@
The remainder of this paper is organized as follows: Section II reviews related work on USV-AUV collaboration in underwater tasks. Section III details the methodology of this work. Section IV describes the environmental simulations and experimental results. Finally, Section V concludes the paper and outlines future directions.
@@ II-A Advances in USV-AUV Collaborative Systems
The USV-AUV collaborative system has garnered significant attention due to its potential in marine exploration and environmental monitoring. Early research primarily addressed communication and coordination challenges, including acoustic communication modems and surface relay systems [18], as well as task allocation mechanisms where USVs serve as mobile base stations [19]. Concurrently, cooperative navigation frameworks developed for GPS-denied environments significantly enhanced AUV positioning capabilities [20].
@@ II-B FIM-based Approaches for USV
For USVs operating in dynamic marine environments, FIM has been applied to adaptive sampling strategies, where vehicle trajectories are optimized to improve environmental field reconstruction (e.g., salinity or temperature mapping) [24]. These approaches often combine FIM metrics with gradient-based or sampling-based planners to balance exploration efficiency with motion constraints [25].
@@
Recent advances have focused on integrating FIM into real-time USV path planning under environmental disturbances and communication constraints. Cooperative scenarios, where USV and multiple AUVs collaboratively maximize joint FIM determinant, have also been explored using distributed optimization techniques [14]. Challenges persist in computational scalability for high-dimensional FIM calculations and in robustifying FIM metrics against sensor noise and model mismatches. Emerging solutions include Gaussian process-based approximations of FIM [26] and hybrid learning-control architectures that adaptively tune information-driven objectives [27].
@@
II-C RL-based AUV Control and Planning
Reinforcement learning (RL) has become a prominent approach for AUV control and planning in complex marine environments. Early studies employed value-based methods like Q-learning for basic navigation tasks [28], while later work adopted policy gradient methods to handle dynamic currents [29]. The introduction of deep RL enabled end-to-end learning from sensor data [30], though these methods faced challenges in partial observability and data efficiency.
Recent advances focus on multi-agent systems and hybrid approaches. Cooperative RL allows AUV teams to collaboratively perform tasks [31], while model-based RL improves sample efficiency using oceanographic models [32]. Current research addresses communication constraints and safety guarantees through hierarchical RL and meta-learning [33], though real-world deployment challenges remain.
@@ III-A System Architecture
Specifically, the system employs FIM to maximize AUV localization precision by minimizing the determinant of the FIM, thereby reducing estimation uncertainty in the underwater environment. The USV dynamically adjusts its path planning based on real-time FIM calculations, ensuring optimal positioning support for the AUVs. This mathematical approach enhances system robustness, even when external disturbances—such as tidal waves or turbulence—degrade sensor measurements.
To further improve adaptability, the system integrates an RL framework where the state space incorporates environment-aware data, and the reward function evaluates USV-AUV collaboration efficiency. This enables multi-AUV systems to learn and refine navigation and task execution strategies under extreme conditions, increasing mission complete efficiency. By continuously adapting to real-time environmental feedback, the RL component ensures sustained performance optimization.
@@ III-B FIM-based USV Path Planning for AUV Positioning
Our system enables accurate underwater vehicle positioning through an optimized USV navigation strategy, leveraging USBL technology, while further using the optimization of FIM’s determinant to realize efficient USV path planning to minimize AUV positioning error. As shown in Fig. 1 and Fig. 2, the system architecture consists of:
  • A surface USV located at coordinates $(x,y,\eta)$ equipped with a USBL transducer array (inter-element spacing $d/2$), while using GPS for self-positioning.
@@
The positioning framework is built upon a measurement model in USBL that correlates observed phase differences with AUV locations. This relationship is mathematically expressed as [34]:
@@
The FIM is derived by calculating the expectation of the second partial derivatives of the log-likelihood function. This quantifies how much information the measurements carry about the unknown USV position $\mathbf{X}$: Under the assumption that $\mathbb{E}[\mathbf{Z}_{k}-\mathbf{h}_{k}(\mathbf{X})]=0$, the FIM is expressed as [35]:
@@
The determinant of $\mathbf{J}_{m}$ serves as a scalar metric for overall positioning accuracy. By simplifying $\det(\mathbf{J}_{m})$. For geometric interpretation and optimization, we isolate two critical geometric factors:
@@
where $A_{k}=\frac{p_{k}^{4}-2S^{2}_{k}p^{2}_{k}}{2S_{k}^{6}}$, with $p_{k}^{2}=x_{k}^{2}+y_{k}^{2}$, while ${\rm sin}\gamma_{k}\!\!=\!\!\frac{z_{k}}{S_{k}}$, and $\varphi_{k}=\arctan(y_{k}/x_{k})$ denotes the azimuth angle. The first term emphasizes depth and slant-dependent observability, while the second term penalizes collinear AUV placements. This directly motivates the optimal formations, which reveals that optimal positioning requires:

### v2

@@ Abstract
This paper develops a novel unmanned surface vehicle (USV)–autonomous underwater vehicle (AUV) collaborative system designed to enhance underwater task performance in extreme sea conditions. The system integrates a dual strategy: (1) high-precision multi-AUV localization enabled by Fisher Information Matrix (FIM)-optimized USV path planning, and (2) a Reinforcement Learning (RL)-based cooperative planning and control framework for multi-AUV task execution. Extensive experimental evaluations in the underwater data collection task demonstrate the system’s operational feasibility, with quantitative results showing significant performance improvements over baseline methods. The proposed system exhibits robust coordination capabilities between USV and AUVs while maintaining stability in extreme sea conditions. To facilitate reproducibility and community advancement, we provide an open-source simulation toolkit available at: https://github.com/360ZMEM/USV-AUV-colab .
@@ I Introduction
Despite recent advances, existing USV-AUV frameworks often lack the intelligence and adaptability required for complex tasks in extreme sea conditions [10]. For instance, while Ultra-Short Baseline (USBL) positioning via USV improves localization, most approaches either fix the USV at static coordinates or employ heuristic path planning [11]. Such rigidity results in inefficient USV-AUV linkage and fails to adapt to dynamic ocean environments, leading to suboptimal positioning and task performance. Further, traditional AUV control strategies (e.g., PID, fuzzy control) rely heavily on precise dynamic models, exhibit poor adaptability to turbulence, and struggle with multi-objective optimization or high-dimensional sensor data [12], which has motivated the application of learning-based methods. Crucially, many studies treat ocean dynamics as negligible noise, which is an unrealistic assumption in extreme conditions [13].
@@
To address these challenges, a dual strategy is essential: an adaptive planning and control policy to enable dynamic decision-making in uncertain environments without relying on rigid models while optimizing multi-AUV collaboration for efficient task execution, and an uncertainty- and information-aware path planning method to improve the accuracy of AUV positioning via USV to enhance the robustness of the USV-AUV system, mitigating localization errors caused by extreme sea conditions.
@@
This motivates the integration of reinforcement learning (RL) and Fisher Information Matrix (FIM) optimization, a synergistic approach that jointly tackles control and localization challenges [14]. On one hand, RL’s model-free adaptive optimization enables AUVs to learn robust policies through environmental interaction, overcoming nonlinearities and time-varying disturbances without precise dynamic models [15, 16]. Multi-Agent RL (MARL) further enhances cooperative planning and control, enabling efficient multi-AUV coordination under communication constraints [17]. On the other hand, FIM-based optimization directly quantifies and minimizes positioning uncertainty via USV path planning to maximize the system’s information content, thereby reducing the Cramér-Rao Bound (CRB) for AUV navigation [14]. By combining RL’s adaptability with FIM’s precision, the proposed system achieves both intelligent decision-making and high-accuracy localization, which is a critical requirement for extreme sea scenarios.
@@
Based on the above analysis, this paper develops a FIM and RL-based USV-AUV collaborative system for underwater tasks in extreme sea conditions to improve mission completion efficiency. The contributions are listed as follows:
@@
  • A novel USV-AUV collaborative system is developed, which integrates a FIM- and RL-based dual strategy to enable dynamic decision-making and uncertainty- and information-aware capabilities to improve mission completion efficiency for underwater tasks.
@@
  • This system achieves accurate positioning of AUVs via USV path planning by minimizing the determinant of the system’s FIM. Based on this, by integrating environment-aware abilities into the state space and employing a shaped reward function for multi-AUV collaboration, we further utilize RL to empower the multi-AUV with intelligence and adaptability for the underwater task.
@@
  • Through comprehensive experiments in the underwater data collection task, our system demonstrates superior feasibility and excellent performance, exhibiting robustness under extreme sea conditions.
@@
The remainder of this paper is organized as follows: Section II reviews related work on USV-AUV collaboration in underwater tasks. Section III details the methodology of this work. Section IV describes the environmental simulations and experimental results. Finally, Section V concludes the paper and outlines future directions. For ease of reference, the main symbols and their corresponding explanations used throughout this paper are summarized in Table I.
TABLE I: Main Symbols and Explanations
Symbols Definition
$\mathbf{p}_{\text{target}}^{k}$ Target SN position for AUV $k$
$\mathbf{Z}_{k}$ USBL measurement for AUV $k$
$\mathbf{h}_{k}(\mathbf{X})$ Measurement model
$\Delta\varphi_{x,k},\Delta\varphi_{y,k}$ Phase differences (x/y)
f,c,df,~c,~d Frequency; sound speed; array spacing
[…]
$\gamma_{k},~\varphi_{k}$ Elevation; azimuth of AUV $k$
$A_{k}$ Geometric weight
$\alpha_{ij}=\varphi_{i}-\varphi_{j}$ Azimuth separation
$S_{0},~\gamma_{0}$ Slant/elevation (symmetric case)
$\chi=\sum_{i<j}\sin^{2}\alpha_{ij}$ Angular-diversity factor
$\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma$ State, action, reward, dynamics, discount
$\pi_{\phi}$ Deterministic policy (actor)
$Q_{\theta_{1}},Q_{\theta_{2}}$ Twin critics (Q-functions)
$\phi^{\prime},~\theta_{i}^{\prime}$ Target networks
$\lambda_{Q},~\lambda_{\pi}$ Critic/actor learning rates
$\tau$ Soft-update rate
$\mathcal{D}$ Replay buffer
$\rho^{\pi}(s)$ State visitation
$s_{t}^{k}$ State of AUV $k$ at $t$
$\Delta\mathbf{p}_{j}^{k}$ Rel. pos. to AUV $j$ (norm.)
$\Delta\mathbf{p}_{\text{target}}^{k}$ Rel. pos. to target SN (norm.)
$\tilde{\mathbf{p}}_{t}^{k}$ Self-position (norm.)
$\rho_{\text{overflow}}$ Data-overflow ratio
$\mathbb{I}_{\text{border}}^{k}$ Border flag
$a_{t}^{k}=[v_{\text{norm}},\theta_{\text{norm}}]^{\top}$ Normalized speed/heading
$d_{k}^{\text{target}}$ AUV–target distance
$d_{kj}$ Inter-AUV distance
$N_{\text{DO}}$ Overflow count
$N_{\text{POI}}$ Total number of the underwater sensor nodes
$\mathbb{I}_{\text{TL}}$ Transmission-success flag
$E_{k}$ Energy of AUV $k$
$u,v$ Wave velocities ($x^{\prime},y^{\prime}$)
$\eta$ Free-surface elevation
$\Delta t,\Delta x,\Delta y$ Time/space steps (FDM)
$V_{x,i,j},V_{y,i,j}$ Turbulent velocity (grid)
$\varpi_{i,j}$ Vorticity (grid)
$(x_{0},y_{0})$ Vortex center
$\Gamma,~\delta$ Vortex intensity; core radius
@@ II-A Advances in USV-AUV Collaborative Systems
The USV-AUV collaborative system has garnered significant attention due to its potential in marine exploration and environmental monitoring. Early research primarily addressed communication and coordination challenges, including acoustic communication modems and surface relay systems [18], as well as task allocation mechanisms in which USVs serve as mobile base stations [19]. Concurrently, cooperative navigation frameworks developed for GPS-denied environments significantly enhance AUV positioning capabilities [20].
@@ II-B FIM-based Approaches for USV
Recent studies have extensively applied the FIM to enhance the efficiency and robustness of USV and AUV operations in dynamic marine environments. The foundational work by Wilson et al. [25] first formulated trajectory optimization via FIM maximization to improve parameter observability, establishing the theoretical groundwork for information-driven mobile robot planning. Building on this, Bahr et al. [7] demonstrated the use of mobile surface beacons for cooperative underwater navigation, aligning with the concept of a dynamic USV acting as a moving information source. Subsequently, Wang et al. [14] integrated USBL measurements with FIM optimization to derive Dubins-path-based USV trajectories that directly improve AUV localization precision, while Paull et al. [18] provided a comprehensive review tracing the evolution from traditional acoustic navigation to information-theoretic trajectory optimization.
@@
In practical applications, FIM has also been widely adopted in adaptive sampling strategies, where USV trajectories are optimized to enhance environmental field reconstruction (e.g., salinity or temperature mapping) [24]. These approaches often combine FIM metrics with gradient-based or sampling-based planners to balance exploration efficiency and motion constraints [25]. More recent efforts have extended FIM to real-time USV path planning under environmental disturbances and communication limitations, as well as to cooperative USV–AUV systems that maximize the joint FIM determinant through distributed optimization [14]. Despite these advances, challenges remain in computational scalability for high-dimensional FIM calculations and in maintaining robustness under sensor noise and model mismatches. Emerging solutions include Gaussian process-based FIM approximations [28] and hybrid learning-control architectures that adaptively tune information-driven objectives [29].
Building upon these theoretical and practical foundations, our work extends FIM-based trajectory optimization into a multi-agent coordination framework, where the FIM explicitly models how the USV’s sensing and motion contribute to AUV localization accuracy in real time. This tight integration of information-theoretic modeling with RL-based control establishes a principled bridge between estimation theory and cooperative autonomy, constituting a key methodological innovation of our system.
II-C RL-based AUV Control and Planning
RL has become a prominent approach for AUV control and planning in complex marine environments. Early studies employed value-based methods such as Q-learning for basic navigation and station-keeping tasks [30], while subsequent work adopted policy gradient algorithms to cope with dynamic ocean currents and uncertain disturbances [31]. With the advent of RL, end-to-end control frameworks emerged that directly learn from raw sensor data [32], greatly improving adaptability but still facing challenges in partial observability, sample inefficiency, and safety assurance. To overcome these limitations, recent research has shifted toward hybrid, model-based, and cooperative RL frameworks that incorporate environmental priors and physical constraints. For example, Zhang et al. [38] introduced a hybrid control architecture combining RL with model-based control to enhance robustness and stability, while Wang et al. [39] integrated Model Predictive Control (MPC) and Control Barrier Functions (CBF) within an RL framework to adaptively tune control weights, achieving reliable trajectory tracking and obstacle avoidance in complex underwater conditions. Similarly, Xi et al. [40] proposed the Comprehensive Ocean Information D3QN (COID), which exploits real oceanographic data and a double dueling deep Q-network to enable flexible and data-driven path planning across dynamic marine environments.
In parallel, meta-learning and hierarchical mechanisms have been introduced to improve sample efficiency and generalization in non-stationary settings, as demonstrated by Fan et al. [41], while multi-agent cooperative RL allows teams of AUVs to coordinate complex missions under communication constraints [33]. Model-based RL approaches have also leveraged ocean simulation and real data assimilation to accelerate convergence and improve safety guarantees [34, 35]. Furthermore, Wozniak et al. [42] applied the twin-delayed deep deterministic policy gradient algorithm to control hydrobatic AUVs, demonstrating robust depth and pitch regulation in simulations and revealing the potential of RL-based controllers for agile underwater maneuvers, despite sim-to-real challenges.
Collectively, these studies demonstrate that RL-based AUV control and planning have evolved from simple navigation and motion control to multi-objective, cooperative, and constraint-aware optimization, forming a solid foundation for adaptive underwater autonomy. However, most existing RL frameworks still focus on single-vehicle control under simplified conditions, lacking treatment of multi-AUV cooperation and localization coupling. To address these gaps, our work integrates information-driven positioning and cooperative RL within a unified framework. Specifically, the USV optimizes its trajectory to maximize the FIM, improving AUV localization accuracy, while the resulting high-fidelity positional information is embedded into the multi-agent RL loop to enhance coordination and situational awareness. Moreover, the USV–AUV localization relationship is mathematically formulated through the FIM, enabling decoupled yet interactive learning between localization and policy optimization. This explicit information-coupled design offers interpretability, modularity, and theoretical rigor, marking a key methodological novelty that bridges information-theoretic modeling with reinforcement learning-based multi-AUV coordination in complex marine environments.
@@ III-A System Architecture
Specifically, the system employs FIM to maximize AUV localization precision by minimizing the determinant of the FIM, thereby reducing estimation uncertainty in the underwater environment. The USV dynamically adjusts its path planning based on real-time FIM calculations, ensuring optimal positioning support for the AUVs. This mathematical approach enhances system robustness, even when external disturbances such as tidal waves or turbulence degrade sensor measurements.
To further improve adaptability, the system integrates an RL framework in which the state space incorporates environment-aware data, and the reward function evaluates the USV-AUV collaboration efficiency. This enables multi-AUV systems to learn and refine navigation and task execution strategies under extreme conditions, thereby increasing mission completion efficiency. By continuously adapting to real-time environmental feedback, the RL component ensures sustained performance optimization.
@@ III-B FIM-based USV Path Planning for AUV Positioning
Our system enables accurate underwater vehicle positioning through an optimized USV navigation strategy that leverages USBL technology while further utilizing the optimization of the FIM’s determinant to achieve efficient USV path planning, thereby minimizing AUV positioning errors. As shown in Fig. 1 and Fig. 2, the system architecture consists of:
  • A surface USV located at coordinates $(x,y,\eta)$ equipped with a USBL transducer array (inter-element spacing $d/2$) while using GPS for self-positioning.
@@
The positioning framework is built upon a measurement model in USBL that correlates observed phase differences with AUV locations. This relationship is mathematically expressed as [USBL]:
@@
The FIM is derived by calculating the expectation of the second partial derivatives of the log-likelihood function. This quantifies how much information the measurements carry about the unknown USV position $\mathbf{X}$: Under the assumption that $\mathbb{E}[\mathbf{Z}_{k}-\mathbf{h}_{k}(\mathbf{X})]=0$, the FIM is expressed as [FIM]:
@@
The determinant of $\mathbf{J}_{m}$ serves as a scalar metric for overall positioning accuracy. By simplifying $\det(\mathbf{J}_{m})$ for geometric interpretation and optimization, we isolate two critical geometric factors:
@@
where $A_{k}=\frac{p_{k}^{4}-2S^{2}_{k}p^{2}_{k}}{2S_{k}^{6}}$, with $p_{k}^{2}=x_{k}^{2}+y_{k}^{2}$, while ${\rm sin}\gamma_{k}\!\!=\!\!\frac{z_{k}}{S_{k}}$, and $\varphi_{k}=\arctan(y_{k}/x_{k})$ denotes the azimuth angle. The first term emphasizes depth and slant-dependent observability, while the second term penalizes collinear AUV placements. This directly motivates the optimal formations, revealing that optimal positioning requires:
@@
For the convenience of research without loss of generality, this work only considers the influence of the relative angles between AUVs and the target on positioning accuracy. When the distance between the AUVs and the USV is equal and the AUVs are at the same depth, only the angle of the AUVs is considered, and the influence of AUV formation on positioning accuracy is only related to ${\rm\!sin}^{2}\alpha_{ij}$:
where ${\rm sin}\gamma_{0}\!\!=\!\!\frac{z_{k}}{S_{0}}$, while $\chi\!\!=\!\!\sum_{1\leq i\textless j\leq m}^{m}{\rm\!sin}^{2}\alpha_{ij}$, with $\alpha_{ij}\!\!=\!\!\varphi_{i}\!-\!\varphi_{j}$ representing the angle between the projections of two AUVs and the USV.
By calculating the waypoints using Eq. (8) with the maximum $\rm det\left(\mathbf{{\textbf{{J}}}}_{m}\right)$, we can implement USV path planning to minimize AUV positioning errors through the above process.
III-C RL-Driven Multi-AUV Decision-Making
Our approach leverages the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm, which extends the Deep Deterministic Policy Gradient (DDPG) algorithm with three key improvements: (1) clipped double Q-learning to mitigate overestimation bias, (2) delayed policy updates to stabilize training, and (3) target policy smoothing to prevent overfitting to narrow peaks in the value function [TD3]. These characteristics make TD3 particularly suitable for underwater vehicle control, where precise continuous actions are required and the dynamics are highly coupled across degrees of freedom.
$\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P}(\cdot|s,a),\gamma>,$ (9)
where continuous state space $\mathcal{S}$, continuous action space $\mathcal{A}$, and reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. $\mathcal{P}(\cdot|s,a)\sim(0,1)$ is the state transition probability. $\gamma\in[0,1]$ represents the discount factor.
Given the state space $\mathcal{S}$ and the action space $\mathcal{A}$, our objective in the underwater tasks is to obtain the expert policy by maximizing the discounted cumulative reward:
where $\pi:\mathcal{S}\to\mathcal{A}$ is a deterministic policy. The agent’s objective is to find the optimal policy $\pi^{*}$ that maximizes the expected return, and $\gamma\in[0,1)$ is the discount factor. This infinite-horizon formulation requires careful treatment of the value function approximation.